OpenStudy (anonymous):

For what rational number "c" do the equations x^3+cx^2+3=0 and x^2+cx+1=0 have c common solution?

myininaya (myininaya):

Try multiplying the second equation by x and adding them together to find what x value they are the same. Then plug in that x to find c.

myininaya (myininaya):

or subtracting

OpenStudy (anonymous):

 (-3)^3+c(-3)^2+3=0 c=25/9 (-3)^3+c(-3)^2-3=0 c=10/3